A hollow conducting sphere has inner radius of a and outer radius of b. If a charge of Q is located at its center and the hollow sphere has a net charge of + 30 . what is the charge distribution in the hollow sphere ?


1
Expert's answer
2020-09-22T15:35:04-0400

"\frac {KQ}{a}+\frac {KQ}{b}=+30"


Thus charge distribution in the hollow sphere will be:


"+{ {30K(\frac {KQ}{a}+ \frac{KQ}{b}})}"